# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@ 3.1.1.1. Stability for $\psi_{i}\equiv 0$
      (3.60) $\dot{\omega}_{2}(\eta_{2},z_{1})$ $=-c_{1}\,\phi_{2}^{2}(\eta_{2})-c_{2}\,\phi_{3}^{2}(z_{1})-\phi_{3}(z_{1})\phi_{1}(z_{2})-\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z_{1})\phi_{1}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z_{1})-\phi_{2}(\eta_{2})\,\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})$
      $+\phi^{2}_{3}(z_{1})+c_{1}\phi_{3}(z_{1})\,\phi_{2}(\eta_{2}).$
@@
      (3.61) $V_{3}(\eta_{2},z_{1},z_{2})=\theta\omega_{2}(\eta_{2},z_{1})+\lambda_{1}(e^{z_{2}}-1-z_{2}).$
@@
      (3.62) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})$ $=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})+\theta\,\phi_{3}^{2}(z_{1})-\theta\phi_{3}(z_{1})\phi_{1}(z_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z_{1})\phi_{1}(z_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z_{1})-\theta\phi_{2}(\eta_{2})\,\phi_{3}(z_{1})$
      $+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})+\theta c_{1}\phi_{3}(z_{1})\,\phi_{2}(\eta_{2})+\phi_{1}(z_{2})\left(-\phi_{2}(\eta_{2})-\dot{z}_{1}+u^{*}-u\right).$
@@
      (3.63) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})$ $=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})+\theta\,\phi_{3}^{2}(z_{1})-\theta\phi_{3}(z_{1})\phi_{1}(z_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z_{1})\phi_{1}(z_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z_{1})-\theta\phi_{2}(\eta_{2})\,\phi_{3}(z_{1})$
      $+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})+\theta c_{1}\phi_{3}(z_{1})\,\phi_{2}(\eta_{2})+\phi_{1}(z_{2})\left((c_{2}-1)\phi_{3}(z_{1})+\phi_{1}(z_{2})+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z_{1})\phi_{1}(z_{2})-(c_{1}+1)\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\right)$
      $+\phi_{1}(z_{2})\left(u^{*}-u\right).$
@@
      (3.64) $\begin{aligned} u&=u^{*}+c_{3o}\phi_{1}(z_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z_{1})+(c_{2}-1-\theta)\phi_{3}(z_{1})+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z_{1})\phi_{1}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-(c_{1}+1)\phi_{2}(\eta_{2})\\ \ &+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{2})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z_{1})\right),\end{aligned}$
@@
      (3.65) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2}).$
    Therefore, global asymptotic stability follows. ∎
@@
$\begin{aligned} u&=u^{*}+c_{3o}\phi_{1}(z_{2})-\theta\frac{1}{\lambda_{3}}\phi^{2}_{3}(z_{1})+(\dfrac{\lambda_{1}}{\lambda_{3}}-1-\theta)\phi_{3}(z_{1})+\frac{1}{\lambda_{3}}\phi_{3}(z_{1})\phi_{1}(z_{2})-\frac{1}{\lambda_{2}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-(\dfrac{\lambda_{3}}{\lambda_{2}}+1)\phi_{2}(\eta_{2})\\ \ &+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{2})}\left(\frac{1}{\lambda_{2}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{1}{\lambda_{2}}\phi^{2}_{2}(\eta_{2})+(\dfrac{\lambda_{3}}{\lambda_{2}}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z_{1})\right)\end{aligned}$
@@
$\dot{V}_{3}(\eta_{2},z_{1},z_{2})=-\theta\dfrac{\lambda_{3}}{\lambda_{2}}\,\phi_{2}^{2}(\eta_{2})-\theta\dfrac{\lambda_{1}}{\lambda_{3}}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2}).$
@@
Since the Lyapunov control function $V_{3}(\eta_{2},z_{1},z_{2})$ is radially unbounded, its level sets $\mathcal{L}_{c}=\left\{\eta\in\mathbb{R}^{3}\;\middle|\;V_{3}(\eta_{2},z_{1},z_{2})\leq c\right\}$ are bounded for any $c>0$. In particular, one can choose $c$ such that these level sets are contained in the positively invariant set (3.45). Moreover, this set $\mathcal{K}$ is well-defined. Let $(\eta_{1},\eta_{2},\eta_{3})=(\varepsilon,0,0)$ with $\varepsilon\neq 0$ small enough. Then $z_{1}=0,\quad z_{2}=\varepsilon,$ so that $\phi_{3}(z_{1})=0$, $\phi_{2}(\eta_{2})=0$, and $\phi_{1}(z_{2})=\phi_{1}(\epsilon)\neq 0$. Consequently, the constraint defining $\mathcal{K}$ yields $u^{*}+c_{3o}\phi_{1}(\epsilon)>0$, hence $(\varepsilon,0,0)\in\mathcal{K}$ and thus $\mathcal{K}\neq\varnothing$. Furthermore, $V_{3}(0,0,\varepsilon)=\lambda_{1}(e^{\varepsilon}-1-\varepsilon)\xrightarrow[\varepsilon\rightarrow 0]{}0.$ By continuity, for any $c>0$ there exists $\varepsilon\neq 0$ small enough such that $V_{3}(0,0,\varepsilon)\leq c$. It follows that $(\varepsilon,0,0)\in\mathcal{L}_{c}\cap\mathcal{K}$, and thus $\mathcal{L}_{c}\cap\mathcal{K}\neq\varnothing$.
@@
Figure 2. Fig. 2. Level sets of $V_{3}(\eta_{2},z_{1},z_{2})$ for two fixed values of $z_{2}.$ Colored contours represent the values of the Lyapunov function $V_{3},$ with hotter colors corresponding to larger values. The gray shaded region indicates the positively invariant set $\mathcal{K}$, defined by the system constraints. For each slice, the largest level set of $V_{3}$ entirely contained within $\mathcal{K}$ is highlighted in red. This contour provides a practical estimate of the region of attraction of the equilibrium under the imposed constraints (positivity of the control and admissible state bounds). The plots show that the boundary of $\mathcal{K}$ is strongly influenced by the choice of $z_{2},$ while the geometry of $V_{3}$ remains convex due to its entropic structure. Increasing $\theta$ or $c_{3o}$ tends to enlarge the invariant domain (more dissipation), whereas increasing $c_{2}$ reduces it along the $z_{1}$-direction through the quadratic term in $\phi_{3}(z_{1})$
Figure 3. Figure 3. Time evolution of the closed-loop system under the Lyapunov-based feedback law. The top panel shows the states $\eta_{1}(t),$ $\eta_{2}(t),$ $\eta_{3}(t),$ all converging towards the origin. The middle panel compares the applied control $u(t)$ with the unsaturated control signal $u_{\text{unsat}}(t)$ and the baseline term $u^{*}.$ The gray shaded areas indicate intervals where the control signal reaches the saturation bounds. The bottom panel illustrates the nonlinear feedback component: the function $\phi_{1}(z_{2}(t))$ (left axis) and its contribution to the feedback $u_{\text{fb}}(t)=-k_{u}\phi_{1}(z_{2}(t))$ (right axis). This decomposition highlights the role of the nonlinear terms in shaping the control action. Overall, the plots confirm that the feedback ensures state convergence while maintaining the control within admissible bounds.
@@
    • When $\psi\equiv 0$ (i.e. $r\equiv 0$), there exists $\kappa>0$ and a Lyapunov function $V_{3}(\eta_{2},z_{1},z_{2})$ in (3.61) which is radially unbounded in $(\eta_{2},z_{1},z_{2})$, such that along the closed-loop trajectories
      (3.70) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})\ \leq\ -\,\kappa\Big{(}\phi_{1}(z_{2})^{2}+\phi_{2}(\eta_{2})^{2}+\phi_{3}(z_{1})^{2}\Big{)}.$
@@
      (3.78) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})(t)\ \leq\ -\tfrac{\kappa}{2}W(\eta,z)\ +\ C\,D^{2}e^{-2\sigma t}\ \leq\ C\,D^{2}e^{-2\sigma t},$
      where $W:=\phi_{1}(z_{2})^{2}+\phi_{2}(\eta_{2})^{2}+\phi_{3}(z_{1})^{2}\geq 0$. Integrating on $[0,t]$ yields
      (3.79) $V_{3}(\eta_{2},z_{1},z_{2})(t)\leq V_{3}(\eta_{2},z_{1},z_{2})(0)+\frac{C\,D^{2}}{2\sigma},$
      so $V_{3}(\eta_{2},z_{1},z_{2})(t)$ is uniformly bounded on $0,\infty)$. Since $V_{3}(\eta_{2},z_{1},z_{2})$ is radially unbounded in $(\eta_{2},z_{1},z_{2})$, the trajectories $(\eta_{2},z_{1},z_{2})$ remain in a compact level set of $V_{3}$; in particular $\eta_{2},z_{1},z_{2}$ are bounded. Hence there exists $o_{i}>0$ such that $|\eta_{i}(t)|\leq o_{i}$ for all $t\geq 0$.

### v2

@@ 3.1.1.1. Stability for $\psi_{i}\equiv 0$
      (3.59) $\dot{\omega}_{2}(\eta_{2},z^{3}_{1})$ $=-c_{1}\,\phi_{2}^{2}(\eta_{2})-c_{2}\,\phi_{3}^{2}(z^{3}_{1})-\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})-\phi_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z^{3}_{1})\phi_{2}(\eta_{2})$
      $+\phi^{2}_{3}(z^{3}_{1})+c_{1}\phi_{3}(z^{3}_{1})\,\phi_{2}(\eta_{2}).$
@@
      (3.60) $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=\theta\omega_{2}(\eta_{2},z^{3}_{1})+\lambda_{1}(e^{z^{3}_{2}}-1-z^{3}_{2}).$
@@
      (3.61) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ $=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})+\theta\,\phi_{3}^{2}(z^{3}_{1})-\theta\phi_{3}(z_{1})\phi_{1}(z^{3}_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})-\theta\phi_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})$
      $+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z^{3}_{1})\phi_{2}(\eta_{2})+\theta c_{1}\phi_{3}(z^{3}_{1})\,\phi_{2}(\eta_{2})+\phi_{1}(z^{3}_{2})\left(-\phi_{2}(\eta_{2})-\dot{z}^{3}_{1}+u^{*}-u\right).$
@@
      (3.62) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ $=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})+\theta\,\phi_{3}^{2}(z^{3}_{1})-\theta\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})-\theta\phi_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})$
      $+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z^{3}_{1})\phi_{2}(\eta_{2})+\theta c_{1}\phi_{3}(z^{3}_{1})\,\phi_{2}(\eta_{2})+\phi_{1}(z^{3}_{2})\left((c_{2}-1)\phi_{3}(z^{3}_{1})+\phi_{1}(z^{3}_{2})+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-(c_{1}+1)\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})\right)$
      $+\phi_{1}(z^{3}_{2})\left(u^{*}-u\right).$
@@
      (3.63) $\begin{aligned} u&=u^{*}+c_{3o}\phi_{1}(z^{3}_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z^{3}_{1})+(c_{2}-1-\theta)\phi_{3}(z^{3}_{1})+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-(c_{1}+1)\phi_{2}(\eta_{2})\\ \ &+\frac{\theta\phi_{3}(z^{3}_{1})}{\phi_{1}(z^{3}_{2})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z^{3}_{1})\right),\end{aligned}$
@@
      (3.64) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2}).$
    Therefore, global asymptotic stability follows.
@@
$\begin{aligned} u&=u^{*}+c_{3o}\phi_{1}(z^{3}_{2})-\theta\frac{1}{\lambda_{3}}\phi^{2}_{3}(z^{3}_{1})+(\dfrac{\lambda_{1}}{\lambda_{3}}-1-\theta)\phi_{3}(z^{3}_{1})+\frac{1}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\frac{1}{\lambda_{2}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-(\dfrac{\lambda_{3}}{\lambda_{2}}+1)\phi_{2}(\eta_{2})\\ \ &+\frac{\theta\phi_{3}(z^{3}_{1})}{\phi_{1}(z^{3}_{2})}\left(\frac{1}{\lambda_{2}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-\frac{1}{\lambda_{2}}\phi^{2}_{2}(\eta_{2})+(\dfrac{\lambda_{3}}{\lambda_{2}}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z^{3}_{1})\right)\end{aligned}$
@@
$\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=-\theta\dfrac{\lambda_{3}}{\lambda_{2}}\,\phi_{2}^{2}(\eta_{2})-\theta\dfrac{\lambda_{1}}{\lambda_{3}}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2}).$
@@
Since the Lyapunov control function $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ is radially unbounded, its level sets $\mathcal{L}_{c}=\left\{\eta\in\mathbb{R}^{3}\;\middle|\;V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})\leq c\right\}$ are bounded for any $c>0$. In particular, one can choose $c$ such that these level sets are contained in the positively invariant set (3.44). Moreover, this set $\mathcal{K}$ is well-defined. Let $(\eta_{1},\eta_{2},\eta_{3})=(\varepsilon,0,0)$ with $\varepsilon\neq 0$ small enough. Then $z^{3}_{1}=0,\quad z^{3}_{2}=\varepsilon,$ so that $\phi_{3}(z^{3}_{1})=0$, $\phi_{2}(\eta_{2})=0$, and $\phi_{1}(z^{3}_{2})=\phi_{1}(\epsilon)\neq 0$. Consequently, the constraint defining $\mathcal{K}$ yields $u^{*}+c_{3o}\phi_{1}(\epsilon)>0$, hence $(\varepsilon,0,0)\in\mathcal{K}$ and thus $\mathcal{K}\neq\varnothing$. Furthermore, $V_{3}(0,0,\varepsilon)=\lambda_{1}(e^{\varepsilon}-1-\varepsilon)\xrightarrow[\varepsilon\rightarrow 0]{}0.$ By continuity, for any $c>0$ there exists $\varepsilon\neq 0$ small enough such that $V_{3}(0,0,\varepsilon)\leq c$. It follows that $(\varepsilon,0,0)\in\mathcal{L}_{c}\cap\mathcal{K}$, and thus $\mathcal{L}_{c}\cap\mathcal{K}\neq\varnothing$.
Remark 3.10.
The control in (3.63) may be rewritten as
(3.65) $\begin{aligned} u&=u^{*}+\frac{1}{\phi_{1}(z^{3}_{2})}\left(c_{3}\phi^{2}_{1}(z^{3}_{2})+\theta c_{1}\phi^{2}_{2}(\eta_{2})+\theta c_{2}\phi^{2}_{3}(z^{3}_{1})\right)\\ &+\frac{\phi_{1}(z^{3}_{2})-\theta\phi_{3}(z^{3}_{1})}{\phi_{1}(z^{3}_{2})}\phi_{1}(\eta_{1})-\phi_{2}(\eta_{2})+\frac{\theta\phi_{3}(z^{3}_{1})-\phi_{1}(z^{3}_{2})-\theta\phi_{2}(\eta_{2})}{\phi_{1}(z^{3}_{2})}\phi_{3}(\eta_{3}),\end{aligned}$
@@
Figure 2. Fig. 2. Level sets of $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ for two fixed values of $z_{2}.$ Colored contours represent the values of the Lyapunov function $V_{3},$ with hotter colors corresponding to larger values. The gray shaded region indicates the positively invariant set $\mathcal{K}$, defined by the system constraints. For each slice, the largest level set of $V_{3}$ entirely contained within $\mathcal{K}$ is highlighted in red. This contour provides a practical estimate of the region of attraction of the equilibrium under the imposed constraints (positivity of the control and admissible state bounds). The plots show that the boundary of $\mathcal{K}$ is strongly influenced by the choice of $z^{3}_{2},$ while the geometry of $V_{3}$ remains convex due to its entropic structure. Increasing $\theta$ or $c_{3o}$ tends to enlarge the invariant domain (more dissipation), whereas increasing $c_{2}$ reduces it along the $z^{3}_{1}$-direction through the quadratic term in $\phi_{3}(z^{3}_{1})$
Figure 3. Figure 3. Time evolution of the closed-loop system under the Lyapunov-based feedback law. The top panel shows the states $\eta_{1}(t),$ $\eta_{2}(t),$ $\eta_{3}(t),$ all converging towards the origin. The middle panel compares the applied control $u(t)$ with the unsaturated control signal $u_{\text{unsat}}(t)$ and the baseline term $u^{*}.$ The gray shaded areas indicate intervals where the control signal reaches the saturation bounds. The bottom panel illustrates the nonlinear feedback component: the function $\phi_{1}(z^{3}_{2}(t))$ (left axis) and its contribution to the feedback $u_{\text{fb}}(t)=-k_{u}\phi_{1}(z^{3}_{2}(t))$ (right axis). This decomposition highlights the role of the nonlinear terms in shaping the control action. Overall, the plots confirm that the feedback ensures state convergence while maintaining the control within admissible bounds.
@@
    • When $\psi\equiv 0$ (i.e. $r\equiv 0$), there exists $\kappa>0$ and a Lyapunov function $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ in (3.60) which is radially unbounded in $(\eta_{2},z^{3}_{1},z^{3}_{2})$, such that along the closed-loop trajectories
      (3.70) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})\ \leq\ -\,\kappa\Big(\phi_{1}(z^{3}_{2})^{2}+\phi_{2}(\eta_{2})^{2}+\phi_{3}(z^{3}_{1})^{2}\Big).$
@@
      (3.78) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})(t)\ \leq\ -\tfrac{\kappa}{2}W(\eta,z)\ +\ C\,D^{2}e^{-2\sigma t}\ \leq\ C\,D^{2}e^{-2\sigma t},$
      where $W:=\phi_{1}(z^{3}_{2})^{2}+\phi_{2}(\eta_{2})^{2}+\phi_{3}(z^{3}_{1})^{2}\geq 0$. Integrating on $[0,t]$ yields
      (3.79) $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})(t)\leq V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})(0)+\frac{C\,D^{2}}{2\sigma},$
      so $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})(t)$ is uniformly bounded on $0,\infty)$. Since $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ is radially unbounded in $(\eta_{2},z^{3}_{1},z^{3}_{2})$, the trajectories $(\eta_{2},z^{3}_{1},z^{3}_{2})$ remain in a compact level set of $V_{3}$; in particular $\eta_{2},z^{3}_{1},z^{3}_{2}$ are bounded. Hence there exists $o_{i}>0$ such that $|\eta_{i}(t)|\leq o_{i}$ for all $t\geq 0$.